OpenStudy (anonymous):

For the graph of f(x)= -3+20x-5x2, what is the x coordinate of its vertex? Pleassee hellppp!!

OpenStudy (anonymous):

have you done derivative yet? That would be extremely easy and conceptually easier to understand. If you haven't, then put that equation into this form with the decreasing order f(x) = ax+ bx + c your vertex would have the x-coordinate at (-b/2a). Good luck! To find y, plug in your numerical x-vertex and solve the function

OpenStudy (anonymous):

ax^2+bx+c* sorry for the typo

OpenStudy (anonymous):

Would you PLEASE be able to do it for me? Pleasee im in a hurry, i have a teacher meeting soon!

OpenStudy (anonymous):

f(x)= -3+20x-5x2 x= 2. Plug in formula with a = -5, b = 20

OpenStudy (anonymous):

where do i put the a and b?

OpenStudy (anonymous):

(-b/2a)= vertex (-20/-10) = 2. x= 2

OpenStudy (anonymous):

Ruka: How do i put the numbers in there? && how do i find them?

OpenStudy (anonymous):

2 x=2

OpenStudy (anonymous):

vertix is (2,17)

OpenStudy (anonymous):

Your former "b" value, in order to became your h value must be divided by 2 and squared.

OpenStudy (anonymous):

ok, so after i do what you put in your previous previous answer, how is it written? Caz i dont know how to multyply what by wat and then how its written.. sorry if its too much..

OpenStudy (anonymous):

I know that the answer it 2 but i need to show my work on how i found that.

OpenStudy (anonymous):

so, this is what i have written down on what you told me so far. f(x)=-3+20x-5x2 f(x)=-5x2+20x-3 -5x2+20x-3 -5(x2-4x)-3 -4/2=-2 = -22 (the second 2 is supposed to be on top of the first 2) =4 -5(x2-4x+4)-3+y -5(x2)2-3-20 -5(x-2)2-23 If i did it wrong can you rewrite all that correctly for me to see?

OpenStudy (anonymous):